Climbing ropes stretch when they catch a falling climber, thus increasing the time it takes the climber to come to rest and reducing the force on the climber. In one standardized test of ropes, an 80 kg mass falls 4.8 m before being caught by a 2.5-m-long rope. If the net force on the mass must be kept below 11 kN, what is the minimum time for the mass to come to rest at the end of the fall?
Answer #1

Acceleration, a = F/m = 11000/80 = 137.5 m/s^2

Initial velocity, u = sqrt(2 x 9.8 x 4.8) = 9.7 m/s

Time, t = u/a = 9.7/137.5 = 0.071 sec

Comment in case any doubt please rate my answer....